# RISK: A Framework for GUI Agents in E-commerce Risk Management

### v1

@@ Abstract
E-commerce risk management requires aggregating diverse, deeply embedded web data through multi-step, stateful interactions, which traditional scraping methods and most existing Graphical User Interface (GUI) agents cannot handle. These agents are typically limited to single-step tasks and lack the ability to manage dynamic, interactive content critical for effective risk assessment. To address this challenge, we introduce RISK, a novel framework designed to build and deploy GUI agents for this domain. RISK integrates three components: (1) RISK-Data, a dataset of 8,492 single-step and 2,386 multi-step interaction trajectories, collected through a high-fidelity browser framework and a meticulous data curation process; (2) RISK-Bench, a benchmark with 802 single-step and 320 multi-step trajectories across three difficulty levels for standardized evaluation; and (3) RISK-R1, a R1-style reinforcement fine-tuning framework considering four aspects: (i) Output Format: Updated format reward to enhance output syntactic correctness and task comprehension, (ii) Single-step Level: Stepwise accuracy reward to provide granular feedback during early training stages, (iii) Multi-step Level: Process reweight to emphasize critical later steps in interaction sequences, and (iv) Task Level: Level reweight to focus on tasks of varying difficulty. Experiments show that RISK-R1 outperforms existing baselines, achieving a 6.8% improvement in offline single-step and an 8.8% improvement in offline multi-step. Moreover, it attains a top task success rate of 70.5% in online evaluation. RISK provides a scalable, domain-specific solution for automating complex web interactions, advancing the state of the art in e-commerce risk management.
Figure 1: RISK framework for GUI agents in e-commerce risk management. Left: Task composition for GUI agents in e-commerce risk management, including information search and website verification tasks. Right: RISK framework, which consists of three key components: RISK-Data, RISK-Bench, and RISK-R1. RISK-R1 achieves SOTA performance in this domain.
@@ 1 Introduction
To harness the full potential of GUI agents in this domain, we propose a novel Web UI agent framework, called RISK, which comprises three key components: (1) RISK-Data. RISK-Data is collected using the Browser Use framework (Müller & Žunič, 2024), which is a framework that integrates advanced context management, optimized prompt templates for both page screenshots and HTML DOM structures, and precise low-level interaction capabilities. We aim to systematically distill and embed the framework’s advanced knowledge into the data, thereby improving the success rate of multi-step, real-world web workflows. After a meticulous curation process, RISK-Data contains 8,492 single-step and 2,386 multi-step interaction trajectories on various task scenarios, shown in Figure 1. (2) RISK-Bench. RISK-Bench is collected for evaluating the performance of GUI agents in e-commerce risk management. It consists of 802 single-step and 320 multi-step trajectories, which are graded into three difficulty levels: easy, moderate, and difficult. (3) RISK-R1. RISK-R1 is an RFT framework based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024). We design a framework-driven reward function and optimization objective to effectively guide the learning process of GUI agents and enable a seamless transition from training to deployment. Specifically, there are four key aspects: (i) Output Format: Updated format reward that enhances the syntactic correctness of the model’s output and task understanding, (ii) Single-step Level: Stepwise accuracy reward that measures action accuracy considering both action completeness and training process, (iii) Multi-step Level: Process reweight that emphasizes the step stage in the interaction process, and (iv) Task Level: Level reweight that focuses on different difficulty levels of tasks.
@@
Experiments on RISK-Bench demonstrate that our approach achieves substantial gains over existing baselines in e-commerce risk management tasks. In offline evaluation, RISK-R1-7B improves single-step performance by 6.8% and multi-step performance by 8.8%. In online evaluation, it attains a top task success rate of 70.5%. Comprehensive analysis further validates the synergistic contributions of each component in RISK-R1. Our contributions are summarized as follows:
@@
  • We introduce the RISK framework, which integrates domain-specific data collection, benchmarking, and reinforcement fine-tuning for GUI agents in e-commerce risk management.
@@
  • We develop RISK-Data, a high-quality dataset with 8,492 single-step and 2,386 multi-step interaction trajectories, and RISK-Bench, a benchmark with 802 single-step and 320 multi-step trajectories for evaluating GUI agents in this domain.
@@
  • We propose RISK-R1, a novel RFT approach based on GRPO, with a comprehensive reward function and optimization objective to enhance the learning process of GUI agents and facilitate deployment in real-world applications.
@@
  • Extensive experiments demonstrate that RISK-R1 outperforms existing baselines, achieving SOTA results in both offline and online evaluations on e-commerce risk management tasks.
@@ 2.1 GUI Agents
GUI agents are intelligent systems that can understand and interact with graphical user interfaces through various actions (e.g., click, type), to accomplish automated execution of complex GUI tasks (Sun et al., 2024; Zhang et al., 2024; Tang et al., 2025; Hu et al., 2025). GUI agents can be broadly categorized into three types: (1) Expert knowledge-driven workflow. These agents construct a workflow consisting of two main components (Li et al., 2024; Wang et al., 2025; Xie et al., 2025; Zhang et al., 2025; Jiang et al., 2025): planner and actioner, where planner decomposes high-level tasks into sub-tasks and generates corresponding action sequences, and actioner is responsible for providing an accurate element localization (e.g., bounding box, DOM tree index). However, these agents heavily rely on expert knowledge to design the workflow and cause error accumulation in long-horizon tasks. (2) Data-driven training. These agents are MLLMs trained on GUI understanding and interaction datasets through supervised fine-tuning (SFT) (Wu et al., 2024; Xu et al., 2024; Qin et al., 2025; Lin et al., 2025) or reinforcement fine-tuning (RFT) (Luo et al., 2025; Zhou et al., 2025; Tang et al., 2025; Liu et al., 2025). Rather than decomposing tasks into sub-tasks, they can end-to-end generate actions based on the current GUI state and task instructions. However, it is challenging to deploy these agents in real-world applications due to the poor generalization ability on unseen webpages and the expensive, high-quality data collection. (3) GUI agent framework. Recently, frameworks such as WebVoyager (He et al., 2024), OpenManus (Liang et al., 2025) and Browser Use (Müller & Žunič, 2024) garner significant attention in the GUI agent community for several reasons: (1) Customized GUI agents by integrating various LLMs and tools, (2) Enhanced context management and more rigorous handling of tool I/O parameters, (3) Capability to interact with real-world webpages and collect complete trajectories for further model training. Given these advantages, a practical approach to domain-specific GUI agents is to use these frameworks to collect domain-specific data, fine-tune MLLMs (supervised or reinforcement), and redeploy the trained models within the frameworks to further enhance their performance for real-world applications.
@@ 2.2 RFT in GUI Agents
Following the release of DeepSeek-R1 (Guo et al., 2025), RFT with rule-based rewards have been widely adopted (Zhang & Zuo, 2025; Feng et al., 2025; Huang et al., 2025). As RFT is anticipated to tackle the problem of poor generalization in SFT, it is also introduced in GUI tasks (Yuan et al., 2025; Liu et al., 2025). Currently, the mainstream methods focus on designing reward functions to guide the learning of grounding capability. For instance, GUI-R1 (Luo et al., 2025) takes the prediction point within the bounding box of the target element as a successful action and assigns a binary reward accordingly. GUI-G1 (Zhou et al., 2025) further considers the relative size of the grounding box and designs a more fine-grained reward function. GUI-G2 (Tang et al., 2025) proposes a Gaussian continuous reward mechanism for a flexible evaluation of grounding accuracy. However, while interacting with real-world webpages, GUI agent frameworks (Müller & Žunič, 2024) do not use (x,y) coordinates for element selection but employ the element index in the DOM tree combined with various tools. Therefore, the gap between the GUI model training and deployment settings makes the existing reward functions inapplicable, and then the model cannot be employed in GUI agent frameworks flexibly for real-world applications. To address this issue, we propose an RFT framework, named RISK-R1, to train GUI agents for e-commerce risk management. RISK-R1 designs a comprehensive reward function and optimization objective to effectively guide the learning process of GUI agents and enable a seamless transition from training to deployment.
@@ 3.1 Task Design
In practical applications in the e-commerce risk management domain, GUI agents are required with the following capabilities: (1) Information Search: GUI agents should be able to efficiently navigate through various webpages and interfaces to locate specific information, such as transaction details, user profiles, and historical data. This involves understanding the structure of webpages, recognizing relevant elements, and executing appropriate actions to retrieve the needed information. (2) Website Verification: GUI agents must be capable of verifying the authenticity and security of websites. This includes checking for secure connections (e.g., URL redirection), validating certificates, and identifying potential phishing or fraudulent sites. The ability to discern trustworthy sources is crucial in mitigating risks associated with online transactions. Developed based on these capabilities, our task composition is shown in Figure 1 and detailed in Appendix A.2.
Figure 2: Data construction process for GUI agents in e-commerce risk management. By leveraging the Qwen-VL-Max, human-defined prompts, and diversified question templates, the Browser Use framework conducts multi-round interactions with webpages to collect raw data. Then, a series of data post-processing steps is applied to ensure the quality of the dataset.
@@ 3.2 Data Construction
We develop a data construction pipeline, as illustrated in Figure 2, to gather high-quality data. The process begins with leveraging the capabilities of Qwen-VL-Max, a powerful vision-language model, to interact with webpages. We further design human-defined prompts and diversified question templates to guide the interactions, where domain-specific knowledge and scenarios are incorporated to ensure the relevance of the collected data. Based on these, the Browser Use framework facilitates multi-round interactions with various webpages, allowing for the collection of raw data that encompasses a wide range of scenarios encountered in e-commerce risk management.
Following data collection, we implement a series of post-processing steps to refine and curate the dataset. This includes (1) Trajectory Filtering: We filter out incomplete or unsuccessful interaction trajectories to ensure the meaningfulness of the data. (2) Step Cleaning: In the successful trajectories, there are some redundant or failed steps (e.g., repeatedly circumventing a slider captcha). We clean these steps to prevent the model from learning incorrect behaviors. (3) Information Refinement: We extract and structure the information (e.g., removing one-shot examples) from the raw data to facilitate easier access and analysis. (4) Data Augmentation: We apply various augmentation techniques to enhance the diversity and robustness of the dataset, such as paraphrasing questions and removing screenshots. (5) Multi-step generation: We generate multi-step interaction sequences by chaining together individual steps. As shown in Figure 2, we replace prompts with the last step’s response after the first step, forming a “think-action-observation” loop and reducing trajectory length. Compared with single-step samples, this helps to simulate more complex scenarios that GUI agents may encounter in real-world applications. (6) Difficulty Grading: We categorize the data into different difficulty levels based on the accuracy of the advanced MLLM’s response. This allows for a curriculum learning in the training process and a more nuanced evaluation of GUI agents’ performance across varying levels of task complexity. Ultimately, we obtain a high-quality dataset and benchmark that effectively supports the development and assessment of GUI agents in this domain.
@@ 3.3 Data Statistics
As shown in Appendix Table 5, RISK-Data comprises 8,492 single-step and 2,386 multi-step interaction trajectories, which are graded into three difficulty levels: easy, moderate, and difficult, based on the accuracy of the advanced MLLM’s response  [^1] We use Qwen-VL-Max to answer each question 5 times. If the accuracy is 100%, 20-80%, and below 20%, we categorize the question as easy, moderate, and difficult, respectively.. In RISK-Data, the easy, moderate, and difficult samples account for 52%, 22%, and 26% in single-step tasks, and 36%, 14%, and 50% in multi-step tasks, respectively. Figure 3 elaborates the action type distribution in RISK-Data, which includes 13 action types. The most frequent action is done, accounting for 27.63%, followed by search_google (26.28%) and click_element_by_index (24.31%). We show action definitions in Appendix Table 6 and 7.
Appendix Figure 8 illustrates the token count and step count distribution of multi-step trajectories in RISK-Data, where we use the token count less than 21000 (around 82.94% of trajectories) for training because of the GPU memory limit. The minimum, maximum, and mean step count of trajectories are 4, 30, and 7.12, respectively. RISK-Bench consists of 802 single-step and 320 multi-step trajectories, where the easy, moderate, and difficult samples account for 47%, 25%, and 28% in single-step tasks, and 30%, 17%, and 53% in multi-step tasks, respectively. To ensure data integrity and prevent leakage, the samples in RISK-Bench are excluded from the training set.
@@ 4.1 Preliminaries
The input at each step of a trajectory consists of the question $q\in Q$, the current webpage screenshot $I_{t}$, and the DOM tree $D_{t}$. The policy model takes these inputs and generates a set of candidate responses $O=\{o_{1},o_{2},\ldots,o_{G}\}$. Each response $o_{i}$ will be evaluated by a reward model to obtain a reward score $r_{i}$. Then group computation is applied on these reward scores to estimate advantages: $A_{i}=\frac{r_{i}-mean(\{r_{1},r_{2},\cdots,r_{G}\})}{std({r_{1},r_{2},\cdots,r_{G}})}$. The policy model is then updated using the optimization objective:
$\mathcal{J}_{\mathrm{GRPO}}(\pi_{\theta})=\mathbb{E}_{q\sim Q,{\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot|q)}}$
$\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\scriptstyle\mathrm{old}}}(o_{i,t}|q,o_{i,<t})}{A}_{i},\mathrm{clip}\Big(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\scriptstyle\mathrm{old}}}(o_{i,t}|q,o_{i,<t})},1\!-\!\epsilon,1\!+\!\epsilon\Big){A}_{i}-\beta\mathrm{D}_{\mathrm{KL}}\left[\pi_{\theta}\parallel\pi_{\mathrm{ref}}\right]\right]\right\},$ (1)
where $\epsilon$ controls the clipping range, $t$ is the token index in the response, and $\beta$ is the coefficient of the KL penalty to constrain the policy from deviating too far from the reference model $\pi_{\mathrm{ref}}$.

### v2

@@ Abstract
E-commerce risk management requires aggregating diverse, deeply embedded web data through multi-step, stateful interactions, which traditional scraping methods and most existing Graphical User Interface (GUI) agents cannot handle. These agents are typically limited to single-step tasks and lack the ability to manage dynamic, interactive content critical for effective risk assessment. To address this challenge, we introduce RISK, a novel framework designed to build and deploy GUI agents for this domain. RISK integrates three components: (1) RISK-Data, a dataset of 8,492 single-step and 2,386 multi-step interaction trajectories, collected through a high-fidelity browser framework and a meticulous data curation process; (2) RISK-Bench, a benchmark with 802 single-step and 320 multi-step trajectories across three difficulty levels for standardized evaluation; and (3) RISK-R1, a R1-style reinforcement fine-tuning framework considering four aspects: (i) Output Format Constraint, (ii) Single-step and (iii) Multi-step Level Reward, and (iv) Task Level Reweight. Experiments show that RISK-R1 achieves a 6.8% improvement in offline single-step and an 8.8% improvement in offline multi-step, using only 7.2% of the parameters of the SOTA baseline. Moreover, it attains a top task success rate of 70.5% in online evaluation. RISK provides a scalable, domain-specific solution for automating complex web interactions in e-commerce risk management. The code is available at https://github.com/RenqiChen/RISK-GUI.
RISK: A Framework for GUI Agents in E-commerce Risk Management
Renqi Chen1,2111This work was done when the first author was an intern at Ant International., Zeyin Tao1, Jianming Guo1, Jingzhe Zhu1, Yiheng Peng1, Qingqing Sun1, Tianyi Zhang1222Corresponding authors., Shuai Chen1222Corresponding authors. 1Ant International, Ant Group 2Fudan University {chenrenqi.crq, zty113091, shuai.cs}@ant-intl.com
Figure 1: RISK framework. Left: Task composition for GUI agents in e-commerce risk management, including information search and website verification tasks. Right: RISK framework, which consists of three key components: RISK-Data, RISK-Bench, and RISK-R1. RISK-R1 achieves SOTA performance in this domain.
@@ 1 Introduction
To harness the full potential of GUI agents in this domain, we propose a novel Web UI agent framework, called RISK, which comprises three key components: (1) RISK-Data. Data is collected using the Browser Use (Müller and Žunič, 2024), which is a framework that integrates advanced context management, optimized prompt templates for both page screenshots and HTML DOM structures, and precise low-level interaction capabilities. We aim to systematically distill and embed the framework’s advanced knowledge into the data, thereby improving the success rate of multi-step, real-world web workflows. After a meticulous curation process, RISK-Data contains 8,492 single-step and 2,386 multi-step interaction trajectories on various task scenarios, shown in Figure 1. (2) RISK-Bench. RISK-Bench is collected for evaluating the performance of GUI agents in e-commerce risk management. It consists of 802 single-step and 320 multi-step trajectories, which are graded into three difficulty levels: easy, moderate, and difficult. (3) RISK-R1. RISK-R1 is an RFT framework based on Group Relative Policy Optimization (GRPO) (Shao et al., 2024). We design a framework-driven reward function and optimization objective to effectively guide the learning process of GUI agents and enable a seamless transition from training to deployment. Specifically, there are four aspects: (i) Output Format: Updated format reward that enhances the syntactic correctness of the model’s output and task understanding, (ii) Single-step Level: Stepwise accuracy reward that measures action accuracy considering both action completeness and training process, (iii) Multi-step Level: Process reweight that emphasizes the step stage in the interaction process, and (iv) Task Level: Level reweight that focuses on different difficulty levels of tasks.
@@
Experiments on RISK-Bench demonstrate that our approach achieves substantial gains over existing baselines in e-commerce risk management tasks. In offline evaluation, RISK-R1-7B improves single-step performance by 6.8% and multi-step performance by 8.8%, using only 7.2% of the parameters of the SOTA baseline. In online evaluation, it attains a top task success rate of 70.5%. Our contributions are summarized as follows:
@@
1) We introduce the RISK framework, which integrates domain-specific data collection, benchmarking, and reinforcement fine-tuning for GUI agents in e-commerce risk management.
@@
2) We develop RISK-Data, a high-quality dataset with 8,492 single-step and 2,386 multi-step interaction trajectories, and RISK-Bench, a benchmark with 802 single-step and 320 multi-step trajectories for GUI agent tasks in this domain.
@@
3) We propose RISK-R1, a novel RFT approach based on GRPO, with a comprehensive reward function and optimization objective to enhance the learning process of GUI agents and facilitate deployment in real-world applications.
@@
4) Extensive experiments demonstrate that RISK-R1 outperforms existing baselines, achieving SOTA results in both offline and online evaluations on e-commerce risk management tasks.
@@ 2.1 GUI Agents
GUI agents are intelligent systems that can understand and interact with graphical user interfaces through various actions (e.g., click, type), to accomplish automated execution of complex GUI tasks (Sun et al., 2024; Zhang et al., 2024; Tang et al., 2025; Hu et al., 2025), which can be broadly categorized into three types: (1) Expert knowledge-driven workflow. These agents construct a workflow consisting of planner and actioner (Li et al., 2024; Wang et al., 2025; Xie et al., 2025; Zhang et al., 2025; Jiang et al., 2025), where planner decomposes high-level tasks into sub-tasks and generates corresponding action sequences, and actioner is responsible for providing an accurate element localization. Agents heavily rely on expert knowledge to design the workflow and cause error accumulation in long-horizon tasks. (2) Data-driven training. Agents are MLLMs trained on GUI understanding and interaction datasets through supervised fine-tuning (SFT) (Wu et al., 2024; Xu et al., 2024; Qin et al., 2025; Lin et al., 2025) or RFT (Luo et al., 2025; Zhou et al., 2025; Tang et al., 2025; Liu et al., 2025). Rather than decomposing tasks into sub-tasks, they can end-to-end generate actions based on the GUI state and task instructions. However, it is challenging to deploy these agents in real-world applications due to the poor generalization ability on unseen webpages and the expensive, high-quality data collection. (3) GUI agent framework. Frameworks such as WebVoyager (He et al., 2024), OpenManus (Liang et al., 2025) and Browser Use (Müller and Žunič, 2024) garner attention for several reasons: 1) Customized GUI agents by integrating various LLMs and tools, 2) Enhanced context management and more rigorous handling of tool I/O parameters, 3) Capability to interact with real-world webpages and collect complete trajectories for further model training. Given these advantages, a practical approach to domain-specific GUI agents is to use these frameworks to collect domain-specific data, fine-tune MLLMs (SFT or RFT), and redeploy the trained models within the frameworks to further enhance their performance for real-world applications.
@@ 2.2 RFT in GUI Agents
Following the release of DeepSeek-R1 (Guo et al., 2025), RFT with rule-based rewards have been widely adopted (Zhang and Zuo, 2025; Feng et al., 2025; Huang et al., 2025). As RFT is anticipated to tackle the problem of poor generalization in SFT, it is also introduced in GUI tasks (Yuan et al., 2025; Liu et al., 2025; Zhou et al., 2025). Currently, the mainstream methods focus on designing reward functions to guide the learning of grounding capability. For instance, GUI-R1 (Luo et al., 2025) takes the prediction point within the bounding box of the target element as a successful action and assigns a binary reward accordingly. GUI-G2 (Tang et al., 2025) proposes a Gaussian continuous reward mechanism for a flexible evaluation of grounding accuracy. However, while interacting with real-world webpages, GUI agent frameworks (Müller and Žunič, 2024) do not use (x,y) coordinates for element selection but employ the element index in the DOM tree combined with various tools. Therefore, the gap between the GUI model training and deployment settings makes the existing reward functions inapplicable, and then the model cannot be employed in GUI agent frameworks flexibly for real-world applications. To address this issue, we propose an RFT framework, named RISK-R1, to train GUI agents for e-commerce risk management. RISK-R1 designs a comprehensive reward function and optimization objective to effectively guide the learning process of GUI agents and enable a seamless transition from training to deployment.
@@ 3.1 Task Design
In practical applications in the e-commerce risk management, GUI agents are required with two capabilities: (1) Information Search: GUI agents should be able to efficiently navigate through various webpages and interfaces to locate specific information, such as transaction details, user profiles, and historical data. This involves understanding the structure of webpages, recognizing relevant elements, and executing appropriate actions to retrieve the needed information. (2) Website Verification: GUI agents must be capable of verifying the authenticity and security of websites. This includes checking for secure connections (e.g., URL redirection), validating certificates, and identifying potential phishing or fraudulent sites. The ability to discern trustworthy sources is crucial in mitigating risks associated with online transactions. Developed based on these capabilities, our task composition is shown in Figure 1 and detailed in Appx. B.
Figure 2: Data construction process. By leveraging the Qwen-VL-Max, prompts, and diversified question templates, the Browser Use framework conducts multi-round interactions with webpages to collect raw data. Then, a series of data post-processing steps is applied to ensure the quality of the dataset.
@@ 3.2 Data Construction
We develop a data construction pipeline, as illustrated in Figure 2, to gather high-quality data. The process begins with leveraging the capabilities of Qwen-VL-Max, a powerful vision-language model, to interact with webpages. We further design prompts and diversified question templates to guide the interactions, where domain-specific knowledge and scenarios are incorporated to ensure the relevance of the collected data. Based on these, the Browser Use framework facilitates multi-round interactions with various webpages, allowing for the collection of raw data that encompasses a wide range of scenarios encountered in e-commerce risk management.
Following data collection, we implement a series of post-processing steps to refine and curate the dataset. This includes (1) Trajectory Filtering: We filter out incomplete or unsuccessful interaction trajectories to ensure the meaningfulness of the data. (2) Step Cleaning: In the successful trajectories, there are some redundant or failed steps (e.g., repeatedly circumventing a slider captcha). We clean these steps to prevent the model from learning incorrect behaviors. (3) Information Refinement: We extract and structure the information (e.g., removing one-shot examples) from the raw data to facilitate easier access and analysis. (4) Data Augmentation: We apply various augmentation techniques to enhance the diversity and robustness of the dataset, such as paraphrasing questions and removing screenshots. (5) Multi-step generation: We generate multi-step interaction sequences by chaining together individual steps. We replace prompts with the last step’s response after the first step, forming a “think-action-observation" loop and reducing trajectory length. Compared with single-step samples, this helps to simulate more complex scenarios that GUI agents may encounter in real-world applications. (6) Difficulty Grading: We categorize the data into different difficulty levels based on the accuracy of the advanced MLLM’s response [^1] We use Qwen-VL-Max to answer each question 5 times. If the accuracy is 100%, 20-80%, and below 20%, we categorize the question as easy, moderate, and difficult, respectively.. This allows for a curriculum learning in the training process and a more nuanced evaluation of GUI agents’ performance across varying levels of task complexity. Ultimately, we obtain a high-quality dataset and benchmark that effectively supports the development and assessment of GUI agents in this domain.
@@ 3.3 Data Statistics
As shown in Appx. Table 6, RISK-Data comprises 8,492 single-step and 2,386 multi-step interaction trajectories, which are graded into three difficulty levels: easy, moderate, and difficult. In RISK-Data, the easy, moderate, and difficult samples account for 52%, 22%, and 26% in single-step tasks, and 36%, 14%, and 50% in multi-step tasks, respectively. Figure 6 elaborates the action type distribution in RISK-Data, which includes 13 action types. The most frequent action is done, accounting for 27.63%, followed by search_google (26.28%) and click_element_by_index (24.31%). We show action definitions in Appx. Table 7 and 8.
Appx. Figure 7 illustrates the token count and step count distribution of multi-step trajectories in RISK-Data, where we use the token count less than 21000 (around 82.94% of trajectories) for training because of the GPU memory limit. The minimum, maximum, and mean step count of trajectories are 4, 30, and 7.12, respectively. RISK-Bench consists of 802 single-step and 320 multi-step trajectories, where the easy, moderate, and difficult samples account for 47%, 25%, and 28% in single-step tasks, and 30%, 17%, and 53% in multi-step tasks, respectively. To prevent data leakage, the samples in RISK-Bench are excluded from the training set.
@@ 4.1 Preliminaries
The input at each step of a trajectory consists of the question, the current webpage screenshot, and the DOM tree. The policy model takes these inputs and generates $G$ candidate responses. Each response will be evaluated by a reward model $R$ for a score. Then group computation is applied on these reward scores to estimate advantages: $A_{i}=\frac{R(\tau_{i})-mean(\{R(\tau_{j})\}^{G}_{j=1})}{std(\{R(\tau_{j})\}^{G}_{j=1})}$. The policy model is then updated using the optimization objective:
$\mathcal{J}(\theta)=$ $\mathbb{E}_{{\tau\sim\pi_{\theta_{\mathrm{old}}}}}\Big[\sum_{t}\min(r_{t}(\theta){A}_{t},\mathrm{clip}(r_{t}(\theta),1\!-\!\epsilon,$
$1\!+\!\epsilon){A}_{t})-\beta\mathrm{D}_{\mathrm{KL}}\left[\pi_{\theta}\parallel\pi_{\mathrm{ref}}\right]\Big],$ (1)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{old}(a_{t}|s_{t})}$ is the probability ratio, $\epsilon$ controls the clipping range, and $\beta$ is the coefficient of the KL penalty to constrain the policy from deviating too far from the reference model $\pi_{\mathrm{ref}}$.